OpenStudy (astrophysics):

Eliminate the parameter to find a Cartesian equation of the curve. Sketch the curve.

OpenStudy (astrophysics):

\[x=\tan^2 \theta,~~~y=\sec \theta, ~~~~ -\pi/2 < \theta < \pi/2\]

OpenStudy (astrophysics):

I'm thinking trig identities...?

OpenStudy (dan815):

mhmm

OpenStudy (dan815):

tan^2+1=sec^2

OpenStudy (astrophysics):

Yeah, so let x = tanx and y = secx?

OpenStudy (astrophysics):

\[1+x^2=y^2 \implies y = \sqrt{1+x^2}\]

OpenStudy (rational):

`sec` takes both positive and negative values

OpenStudy (astrophysics):

But the interval only allows us to take positive square roots, yeah?

OpenStudy (rational):

yeah i think you're right but im not sure if we need to preserve the range of parameter after switching to cartesian form

OpenStudy (rational):

For example : x = t, y = t^2, 0<t<1

OpenStudy (rational):

eliminating parameter gives you y = x^2 do you worry about the range of "t" after switching to cartesian ?

OpenStudy (rational):

we need to loose that info in cartesian form y = sqrt(1+x^2) looks wrong to me because somehow you're trying to forcibly use the range of values of t hmm

OpenStudy (astrophysics):

\[1+x=y^2 \implies y = \sqrt{1+x},~~ x \ge 0\]

OpenStudy (rational):

Another example \(x = r\cos \theta \), y = \(r\sin \theta\) ; \(\large 0\le \theta \le \pi\) Is the cartesian form \(\large x^2 + y^2 = r^2\) or \(\large y = \sqrt{r^2-x^2}\) ?

OpenStudy (rational):

Ah that makes sense so you want to define the same curve by restricting the domain

OpenStudy (rational):

i need to review the general convention, but you make more sense :)

OpenStudy (astrophysics):

Yeah, haha, I'm still kind of confused by this, maybe it's just this question is like the hardest one from the book, and I haven't done enough practice.

OpenStudy (rational):

you're right, start and endpoints matter in cartesian form too even through we loose the orientation of curve information http://tutorial.math.lamar.edu/Classes/CalcII/ParametricEqn.aspx

OpenStudy (astrophysics):

Thanks, for the clarification :)

OpenStudy (rational):

:D